40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 12 01:51:23 2008, rob, Update, Locking, report lock_control_sequence.png
    Reply  Thu Jun 12 15:09:33 2008, alan, Update, Locking, report 
       Reply  Thu Jun 12 15:55:15 2008, rob, Update, Locking, report 
Message ID: 531     Entry time: Thu Jun 12 01:51:23 2008     Reply to this: 532
Author: rob 
Type: Update 
Category: Locking 
Subject: report 
rob, john

We've been working (nights) on getting the IFO locked this week. There's been fairly steady incremental progress each night, and tonight we managed to control CARM(MCL) using PO-DC, with the CARM(AO) path also on PO-DC. In the past, reaching this state has usually meant we're home free, as we could just crank the gain on the common mode servo and merrily reduce the CARM offset. Tonight, however, this state has been very twitchy, and efforts to ramp up the gain have been unsuccessful.

I've attached a diagram which I hope makes clear where we are in the stages of lock acquisition.
Attachment 1: lock_control_sequence.png  77 kB  Uploaded Thu Jun 12 03:40:44 2008  | Hide | Hide all
lock_control_sequence.png
ELOG V3.1.3-